OpenStudy (anonymous):

I need help with this equation, I'm partially done - but I can't figure it out completely.. 6 = x + 2 / 3

OpenStudy (anonymous):

you need to find x?

OpenStudy (anonymous):

is that 6= (x+2)/3 or 6 = x +2/3

OpenStudy (anonymous):

its easy think about it this way if ur trying to get 6 and the exuation u get is x+2 / 3 then u will kind of do the opposite so u will do 6 times three so the equation is x+2= 18 then uminus 2 fromboth sides to get x= 16 then u check it and substitute x with 16 so the equestion is now 16+2 /3 =6 then simplify further 18/3=6 then finish 6=6 so the answer is that x = 16 plzzz give me a medal

OpenStudy (anonymous):

Thank you. c:

OpenStudy (anonymous):

It is the first option, Kashan.

OpenStudy (anonymous):

I've found my answer.

OpenStudy (anonymous):

Ok thats great.. :)